OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 4x - 5y = 24? Choose all answers that are correct. A. (9, 12) B. (1, 4) C. (4, 8) D. (6, 0)

OpenStudy (anonymous):

@Clementine @DarkMoonZ

OpenStudy (anonymous):

@ReadytoLearn_Student

OpenStudy (anonymous):

Hey William! Not the greatest are this...maybe @iGreen can help :)

OpenStudy (anonymous):

ok(: @iGreen?

OpenStudy (anonymous):

@Tazmaniadevil

OpenStudy (anonymous):

@courtney-ray

OpenStudy (anonymous):

Hmmm....@satellite73

OpenStudy (anonymous):

4x+5y=24 Simple and best practice solution for 4x+5y=24 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 4x+5y=24 equation: Simplifying 4x + 5y = 24 Solving 4x + 5y = 24 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5y' to each side of the equation. 4x + 5y + -5y = 24 + -5y Combine like terms: 5y + -5y = 0 4x + 0 = 24 + -5y 4x = 24 + -5y Divide each side by '4'. x = 6 + -1.25y Simplifying x = 6 + -1.25y

OpenStudy (anonymous):

Plugin the point into the equation and see if it is true. For instance Lets do A A: (9,12) 4x - 5y =24 4(9) - 5(12) = 24 36 - 60 = 24 -24 = 24 IS this true or false? If it is true it is an answer to our equation. If it is false it is not an answer

OpenStudy (anonymous):

Oh, high five for great explanations!

OpenStudy (anonymous):

me?